1.) The perimeter of a triangle is 76 centimeters. The second side is twice as long as the first side. The third side is four centimeters shorter than the second side. How long is each side?
Found 2 solutions by haileytucki, mananth:
Answer by haileytucki(390)

76=x+2x+x-4
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
x+2x+x-4=76
Since x and 2x are like terms, add 2x to x to get 3x.
3x+x-4=76
Since 3x and x are like terms, add x to 3x to get 4x.
4x-4=76
Since -4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides.
4x=4+76
Add 76 to 4 to get 80.
4x=80
Divide each term in the equation by 4.
(4x)/(4)=(80)/(4)
Simplify the left-hand side of the equation by canceling the common factors.
x=(80)/(4)
Simplify the right-hand side of the equation by simplifying each term.
x=20
one side is 20
second side is 40
and the third side is 16
20 + 40 +16= 76

Answer by mananth(16949)

let first side be x
second side= = 2x
third side= 2x-4
...
x+2x+2x-4=76
5x-4=76
5x=76+4
5x=80
x=16 cm first side
second side = 2*16= 32 cm
third side = 2x-4 = 2*16 -4 =28 cm